# No Nowhere Constant Continuous Function Maps All Non-Normal Numbers to Normal Numbers

### v3

@@ Abstract
In this work, we consider the set of non-normal numbers $\mathcal{N}^{c}$ and ask if there is a non-empty open interval $I$ and a nowhere constant continuous function $\varphi:I\to\mathbb{R}$ which maps all non-normal numbers to normal numbers, i.e., $\varphi(I\cap\mathcal{N}^{c})\subset\mathcal{N}.$ We answer this question negatively. We call a set with this property super-dense. The super-density of $\mathcal{N}^{c}$ illustrates further its topological size or geometric richness. Surprisingly, the measure-theoretically “larger” set of normal numbers $\mathcal{N}$ does not share this property and we give an explicit Cantor-type construction of a nowhere constant continuous function $\hat{C}$, which maps all normal numbers to non-normal numbers. Our final result discusses the relation between super-density and the Baire categories employing various set-theoretic strategies.
Keywords— Normal numbers, Hausdorff dimension, Cantor function
@@ 1 Introduction and Main Result
Ever since Émile Borel introduced the concept of normality and showed that almost every number is normal [7], generations of mathematicians have worked hard to unravel the secrets of normal numbers. Although many profound statements about the set of normal numbers have been shown [9, 8, 13, 18, 20, 19, 15] and in recent years there has been increasing interest in the computability of concrete normal numbers [1, 4, 5], the most important conjectures concerning normal numbers remain unsolved. In particular, it currently seems unfeasible to show that important mathematical constants like $\sqrt{2},e,$ or $\pi$ are normal.
In contrast, the field of transcendental number theory has made significant breakthroughs. Although many conjectures such as Schanuel’s conjecture regarding transcendental numbers remain open, Hermite, Lindemann and Weierstrass proved in their seminal works the transcendence of $e$ and $\pi$ over a century ago [11, 12, 14, 22]. This disparity raises a natural question: why is it considerably harder to prove a number is normal than to prove it is transcendental? If one understands a proof of normality as a demonstration that a number is not non-normal, the answer seems to be that the set of non-normal numbers is of Hausdorff dimension 1, while the non-transcendental numbers, i.e. the algebraic numbers, are only a countable set. This note presents another perspective on the complexity of non-normal numbers: There is no nowhere constant continuous function that maps all non-normal numbers to normal numbers. This illustrates the intricate geometry of the set of non-normal numbers, potentially explaining the difficulty of proving normality.
To set the stage, we recall first some standard terminology. A real number $x\in{\mathbb{R}}$ can be uniquely expanded with respect to an integer base $b\geq 2$ in the form $x=X_{1}X_{2}\ldots X_{K}.x_{1}x_{2}\cdots$ with digits $X_{i},x_{i}\in\{0,1,\ldots,b-1\}$. This $b$-ary expansion of $x$ is unique if we agree to use a finite expansion, i.e., an expansion with only finitely many nonzero digits, whenever possible. We denote by $\Lambda_{b,d,M}(x)$ the relative fraction of the digit $d$ among the first $M$ fractional $x_{i}$ in base $b$ (see also (2.1)). A real number $x$ is called simply $\mathit{b}$-normal if $\lim_{M\to\infty}\Lambda_{b,d,M}(x)=\frac{1}{b}$, that is if all digits are asymptotically equidistributed in the $b$-ary expansion of $x$. If not just the digits but all finite words $w\in\bigcup_{k=1}^{\infty}\{0,1,\ldots,b-1\}^{k}$ appear with asymptotic frequency $b^{-|w|}$ in the $b$-ary expansion of $x$, we call $x$ a $\mathit{b}$-normal number. Equivalently, a real number $x$ is $b$-normal if and only if $x$ is simultaneously simply $b^{N}$-normal for all natural numbers $N\geq 1$. Finally, $x$ is an absolutely normal number or just a normal number if $x$ is $b$-normal with respect to all integer bases $b\geq 2$. Absolute normality is equivalent to the seemingly weaker condition that a number $x$ is simultaneously simply $b$-normal for all bases $b\geq 2$.
We denote by $\mathcal{N}_{b,s}=\{x\in{\mathbb{R}}\,|\,x\text{ is simply }b\text{-normal}\}$ the set of all $b$-simply normal number for some fixed integer base $b\geq 2$. Analogously, we introduce the set of all $b$-normal real numbers $\mathcal{N}_{b}$, and the set of all (absolutely) normal numbers $\mathcal{N}$. Finally, the complement of a set $A\subset{\mathbb{R}}$ is denoted by $A^{c}:={\mathbb{R}}\setminus A$.
Our results also need the concept of nowhere constant functions:
@@
Note that for a function $\varphi$ to be nowhere constant on some interval $I$, it suffices that the set
@@
is dense in $I$. In particular, all typical realizations of Brownian motions are nowhere constant continuous functions.
Our first main result reads as follows:
Theorem 1.2.
Let $I\subset{\mathbb{R}}$ be any non-empty open interval and $\varphi\colon I\to{\mathbb{R}}$ a nowhere constant continuous function. Then, for every base $b\geq 2$
$\varphi(I\cap\mathcal{N}_{b,s}^{c})\cap\mathcal{N}_{b,s}^{c}\neq\emptyset.$ (1.2)
In fact, we have more generally for every countable collection $(\varphi_{k})_{k\in{\mathbb{N}}}$ of nowhere continuous functions $\varphi_{k}\colon I\to{\mathbb{R}}$,
$\bigcap_{k=1}^{\infty}\varphi_{k}(I\cap\mathcal{N}_{b,s}^{c})\neq\emptyset.$ (1.3)
Of course, (1.2) directly yields
$\varphi(I\cap\mathcal{N}^{c})\cap\mathcal{N}^{c}\neq\emptyset,$
which motivates the title of this note. A few remarks are in order.
  1. 1.
    The proof of Theorem 1.2 is spelled out in Section 2.1 for a single function $\varphi$ and its full generality in Section 2.3. Our method is constructive in the sense that under a few regularity assumptions on the $\varphi_{k}$ we can infer from the proof an algorithm computing inductively the digits of a real number $x$ such that all numbers $\varphi_{k}(x)$ are simultaneously non-normal. The details are presented in Corollary 2.9 below. Theorem 1.2 gives a (nearly) optimal result on joint non-normality, which complements the recent work [15], where the construction of simultaneously normal numbers is addressed. In the future, we hope to address simultaneous normality and non-normality, that is, to construct for example a normal number $x$ such that $x^{2}$ is not normal - mimicking the conjectured property of $\sqrt{2}$.
  2. 2.
    Let us denote by $\bar{\mathbb{Q}}$ the set of real algebraic numbers. Applying Theorem 1.2 to the countable collection $(e^{\alpha x})_{\alpha\in\bar{\mathbb{Q}}\setminus\{0\}}$ and to all open intervals $I=(\alpha,\beta)\subset{\mathbb{R}}$, one concludes that the set of numbers $x$ and all $e^{\alpha x}$ are not normal is a dense subset of $\mathbb{R}$}. Note the remarkable difference compared to the case of algebraicity: according to the celebrated Lindemann-Weierstrass Theorem $e^{x}$ is transcendental for every algebraic $x\neq 0$. While the existence of such numbers is also a consequence of the set of non-normal numbers being of second category [2, 16, 17], as far as we know there has been no constructive method to generate a non-normal number $x\neq 0$ for which $e^{x}$ is not normal. Our proof, however, can easily be turned to an algorithm for the given function $e^{x}$.
  3. 3.
    In the setting of a single nowhere constant function $\varphi$, (1.2) can be strengthened further:
    $\varphi\left(I\cap\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\cap\left(\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\neq\emptyset.$ (1.4)
    This result can be derived by an adaption of our proof strategy of Theorem 1.2 and can be found below as Corollary 2.10.
  4. 4.
    Equation (1.2) might even hold under the weaker condition that $\varphi$ is just a continuous non-constant function. However, our proof fails in this situation and it is unclear to us if the statement of Theorem 1.4 below still holds for non-constant (but not necessarily nowhere constant) functions. We hope to address this question in future research.
In our second main result, we want to argue that Theorem 1.2 should not be seen as a reflection of the measure-theoretic size of $\mathcal{N}^{c}$, but rather as a facet of the topological size and the complicated geometry of $\mathcal{N}^{c}$. To this end, it is convenient to introduce the following concept.
@@
The concept of super-density appears to be novel. It is very vaguely related to the definition of a universally meager set, which is a subset $A\subset X$ of a topological space $X$ such that for every Baire space $Y$ and every nowhere constant continuous map $\varphi:Y\to X$ the preimage $\varphi^{-1}(A)$ is meager in $Y$. [21]. Clearly, a super-dense set is dense in the usual sense with respect to the Euclidean metric; and it is not hard to see that a super-dense set is necessarily uncountable. More will be said in Theorem 1.5 below. Before, we state our second result on (non-)normal numbers.
@@
For every integer $b\geq 2$, there exists a super-dense set $\mathcal{Z}_{b}\subset\mathcal{N}_{b,s}^{c}$ of Hausdorff dimension zero. On the other hand, the set of absolutely normal numbers $\mathcal{N}$ is not super-dense.
The proof of Theorem 1.4 is presented in Section 2.1, where we also give an explicit definition of $\mathcal{Z}_{b}$ in (2.2). We note that Theorem 1.4 implies (1.2) (i.e. Theorem 1.2 for a single function). Due to Borel’s theorem, it might come as a surprise that the non-normal numbers are in some sense “denser” or “richer” than the absolutely normal numbers $\mathcal{N}$, which form a set of full measure. Our interpretation is that the geometry of $\mathcal{N}$ is simpler as $\mathcal{N}$ only contains numbers which behave generically, whereas $\mathcal{N}^{c}$ encodes many fractal-like subsets.
A further intuition for Theorem 1.4 can be found in the fact that $\mathcal{N}$ is a meager set or of first category [2, 16, 17]. In the following theorem we clarify the relation between super-density and Baire categories.
@@
    Let $A\subset\mathbb{R}$ be super-dense. Then, $A$ is also non-meager.
@@
    Let $A\subset\mathbb{R}$ be a co-meager $G_{\delta}$-set, i.e., a countable intersection of open sets. Then, $A$ is super-dense.
@@
The proof of Theorem 1.5 is spelled out in Section 2.2. Note that the first two assertions of Theorem 1.5 directly yield Theorem 1.4 (see also the discussion after Proposition 2.5). This more elegant approach has the disadvantage that it is not constructive and does not give an example of a nowhere constant function $\varphi$ such that $\varphi(\mathcal{N})\cap\mathcal{N}=\emptyset$. We further observe that due to Baire’s category theorem at least one of $A,A^{c}$ as in the third assertion must be non-meager, dense, but not super-dense. Of course, $\mathcal{N}$ is an example for a set which is not super-dense (and meager), and the non-normal numbers $\mathcal{N}^{c}$ are super-dense and non-meager (in fact co-meager). Finally, the sets $A,A^{c}$ from 4. are super-dense, non-meager, but not co-meager.
@@ 2 Proof of Main Results
For the proof of Theorem 1.4, some further notation is useful. Let $b\geq 2$ and $x=0.x_{1}x_{2}x_{3}\ldots$ be the $b$-ary expansion of some $x\in[0,1)$. For $d\in\{0,1,\ldots,b-1\}$ and $M\in{\mathbb{N}}$, we set
$\Lambda_{b,d,M}(x):=\frac{1}{M}\sum_{j=1}^{M}\chi_{d}(x_{j})$ (2.1)
with the characteristic function $\chi_{d}$, that is, $\chi_{d}(d)=1$ and $\chi_{d}(k)=0$ for $k\neq d$. In other words, $\Lambda_{b,M}(x)$ is the density of the digit $d$ in the first $M$ digits of the $b$-ary expansion of $x$. For general $x\in{\mathbb{R}}$, we simply set $\Lambda_{b,d,M}(x):=\Lambda_{b,d,M}(|x|\,\mathrm{mod}\,1).$ We will be interested in the set $\mathcal{Z}_{b}$ of numbers with accumulations of zeros, that is,
We demonstrate first that the Hausdorff dimension $\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0$ as claimed in Theorem 1.4.
Lemma 2.1.
Let $b\geq 2$ be an integer. Then,
$\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0.$ (2.3)
The proof of Lemma 2.1 relies on a standard combinatorial estimate and basic properties of Hausdorff measures. As it does not directly follow from Egglestone’s result on the Hausdorff-dimension of sets with specific asymptotic digit frequencies [10], we choose to sketch the complete argument in the appendix.
Proposition 2.2 below demonstrates that $\mathcal{Z}_{b}$ is super-dense and, thus, establishes the first part of Theorem 1.4.
Proposition 2.2.
For every base $b\geq 2$, the set $\mathcal{Z}_{b}$ is super-dense, that is, for every non-empty interval $I\subset{\mathbb{R}}$ and every nowhere constant continuous function $\varphi\colon I\to{\mathbb{R}}$,
$\varphi(I\cap\mathcal{Z}_{b})\cap\mathcal{Z}_{b}\neq\emptyset.$ (2.4)
Proof.
of numbers with finite $b$-ary expansion and we set
$\tau(x):=\inf\{L\in{\mathbb{N}}\,|\,x=0.x_{1}\ldots x_{L}\}$ (2.6)
for $x\in T$ and otherwise $\tau(x)=\infty$. Our construction a second set
where the inner union runs over all integers $K>\tau(x)^{2}+1$. A number $y\in T^{\prime}$ in the form $x+b^{-K}$ with $x\in T$ and $K>\tau(x)^{2}+1$ and this decomposition is in fact unique as $K$ is determined by the position of the last non-zero digit (which must an isolated ’1’). For a $y=x+\frac{1}{b^{K}}\in T^{\prime}$, we similarly set $\tau^{\prime}(y):=K.$ The function $\tau^{\prime}(y)$ is well-defined by our previous observation. One easily sees that both, $T$ and $T^{\prime}$, are dense sets in $(0,1)$. The idea behind the definition of $T^{\prime}$ is that the zero digits of some $y=x+\frac{1}{b^{K}}\in T^{\prime}$ between the positions $\tau(x)$ and $K$ are preserved under small enough subtractions.
We now construct 4 sequences
Step 1: Choice of starting points and intervals
To construct the two other starting values, we recall that $\varphi$ is nowhere constant, and thus we find some $z\in I_{1,x}$ such that $\varphi(z)\neq\varphi(x_{1}^{(1)}).$ Employing the intermediate value theorem and the density of $T^{\prime}$, we in fact find some $x_{1}^{(2)}\in I_{1,x}$ such that $y_{1}^{(2)}:=\varphi(x_{1}^{(2)})\in T^{\prime}$. Note that the digits $x_{1}^{(2)}$ at position $L_{x,1}$ to $L_{x,1}^{2}$-th are zero.
Step 2: Iterative procedure:
$I_{M+1,x}:=(x_{M+1}^{(1)},x_{M+1}^{(1)}+b^{-L^{2}_{x,M+1}-1})\cap I_{M,x}\neq\emptyset$
[…]
$\Lambda_{b,0,L_{M,x}^{2}}(x)\geq 1-\frac{1}{L_{M,x}},$
It remains to show that the set of absolutely normal numbers $\mathcal{N}$ is not super-dense. In fact, we will construct an even larger set $\mathcal{A}\supset\mathcal{N}$, which is not super-dense. We define
[…]
and recall that the set of $2$-simply normal numbers is given by
We set
$\mathcal{A}:=\mathcal{T}_{1}\cap\mathcal{N}_{s,2}.$ (2.8)
Proposition 2.3.
The set $\mathcal{A}$ is not super-dense.
Proof.
Note that $C$ maps all numbers in $\mathcal{T}_{1}$ to values with finite binary expansion. That is,
$C(\mathcal{A}\cap(0,1))\cap\mathcal{A}=\emptyset.$ (2.9)
To this end, it is convenient to work with the symmetrized version $C_{s}$
which still is continuous and preserves the crucial property of the Cantor function that it maps all numbers with at least a ’1’ in the ternary expansion to a number with finite binary expansion. We now define the enriched Cantor function $\hat{C}:[0,1]\to[0,1]$. To this end, we distinguish the cases, where $x$ has finitely many ’1’s and infinitely many. If $x$ has no ’1’, we set $\hat{C}(x)=C_{s}(x).$ If $x$ has $m$ ’1’s in its ternary expansion, we write $x=0.w_{1}w_{2}\ldots w_{m}\sigma,$ where the finite words $w_{i}$ contain exactly an ’1’ as their last digit - we call such words one-stripped in this proof - and $\sigma$ is a word without any ’1’. Then, we set
$\hat{C}(x):=\left(\sum_{r=1}^{m}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r})\right)+2^{-(m+1)\sum_{k=1}^{r}|w_{r}|}C_{s}(\sigma),$ (2.11)
$\hat{C}(x):=\sum_{r=1}^{\infty}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r}).$ (2.12)
The function $\hat{C}(x)$ is nowhere constant as $C(x)\neq C(y)$ for any $y$ with finite ternary expansion and any $x$ with infinitely many ’0’s and ’2’s.
Next, we verify that $\hat{C}(x)$ is continuous. We consider the approximations $\hat{C}_{k}(x),$ which roughly speaking stop after the $k$-th ’1’ (similarly to the original Cantor function which stops after the first ’1’). A rigorous construction goes as follows. Let $T_{k}:[0,1)\to[0,1)$ be the truncation map
[…]
where the $w_{i}$ are the unique one-stripped words from above with a single ’1’ as their last digit. We simply set $\hat{C}_{k}(x):=\hat{C}(T_{k}(x)).$ Let us show that $\hat{C}_{k}(x)$ is continuous. We note first that $\hat{C}_{1}(x)=C_{s}(x)$ and for each natural number $k$ the difference $\hat{C}_{k+1}-\hat{C}_{k}$ is nonzero only on closed intervals of the form
with one-stripped $w_{i}$ and on each such interval we have
$\hat{C}_{k+1}(0.w_{1}\ldots w_{k}\tau)-\hat{C}_{k+1}(0.w_{1}\ldots w_{k}\tau)=2^{-(k+1)\sum_{j=1}^{r}|w_{j}|}C_{s}(\tau)$
for any ternary word $\tau$. The continuity is evident as one inserts a damped copy of the original (symmetrized) Cantor function in each of the separated intervals $I_{w_{1},w_{2},\ldots,w_{k}}$ , where $\hat{C}_{k}$ is constant. Moreover, by construction
$\sup_{x\in(0,1)}|\hat{C}(x)-\hat{C}_{k}(x)|\leq 2^{-k^{2}}$
so that $\hat{C}_{k}$ converges uniformly to $\hat{C}$. This implies that $\hat{C}$ is continuous, too.
It remains to show that $\hat{C}(\mathcal{T}_{1})\cap\mathcal{N}_{2,s}=\emptyset.$ To this end, let $x\in\mathcal{T}_{1}$ and we write for its ternary expansion $x=0.w_{1}w_{2}\cdots$ with one-stripped $w_{i}$. Let $\hat{w_{i}}$ be the “Cantor image” of $w_{i}$ that is the binary string obtained by flipping each ’2’ in $w_{i}$ to a ’1’. Then, the binary expansion of $\hat{C}(x)$ has the form $\hat{C}(x)=\hat{w_{1}}z_{1}\hat{w_{2}}z_{2}\cdots$, where each $z_{r}$ is a string of zeros. Now, note that after the last digit of $z_{m}$ the binary expansion of insertion of $\hat{C}(x)$ has at most $\sum_{r=1}^{m}w_{r}$ ’1’s but at least $(m-1)\sum_{r=1}^{m}w_{r}$ zeros. That is there a sequence of integers $L_{m}$ s.t.
$\Lambda_{2,0,L_{m}}(C(x))\geq 1-\frac{1}{m}.$
This finishes the proof. ∎
2.2 Proof of Theorem 1.5
@@
In this section, we prove all statements of Theorem 1.5 one by one. The proofs are all independent of each other and are therefore presented separately. We start with the first statement.

### v4

# Super-Dense Sets and Their Role in the Theory of Normal Numbers

@@ Abstract
We introduce and study a new topological notion of the size for subsets of the real line, called super-density. A set $A\subset\mathbb{R}$ is super-dense if for every non-empty open interval $I$ and every nowhere constant continuous function $\varphi\colon I\to\mathbb{R}$, we have $\varphi(I\cap A)\cap A\neq\emptyset$.
We first establish basic properties of super-dense sets. Our main topological result characterizes them within the framework of Baire category: a set with the Baire property is super-dense if and only if it is co-meager.
We then investigate the implications for the theory of normal numbers. We prove that the set of non-normal numbers is super-dense, whereas the set of normal numbers is not. Consequently, no nowhere constant continuous function can map all non-normal numbers to normal numbers. Conversely, we explicitly construct a computable nowhere constant continuous function that maps all normal numbers to non-normal numbers.
Finally, we provide a constructive algorithm that, given any countable family of nowhere constant continuous functions, produces a real number $x$ such that $x$ and all its images under these functions are non-normal. As a corollary, we obtain the existence of a non-normal number $x$ such that $e^{\alpha x}$ is non-normal for every non-zero algebraic $\alpha$.
Keywords— Normal numbers, Baire categories, Baire sets, Super-density, Hausdorff dimension, Cantor function
@@ 1 Introduction and Main Result
Ever since Émile Borel introduced the concept of normality and showed that almost every number is normal [7], generations of mathematicians have worked hard to unravel the secrets of normal numbers. Although many profound statements about the set of normal numbers have been shown [9, 8, 13, 18, 20, 19, 15] and in recent years there has been increasing interest in the computability of concrete normal numbers [1, 4, 5], the most important conjectures concerning normal numbers remain unsolved. In particular, it currently seems infeasible to show that important mathematical constants like $\sqrt{2},e,$ or $\pi$ are normal. By contrast, transcendental number theory has witnessed significant breakthroughs. Although many conjectures such as Schanuel’s conjecture regarding transcendental numbers remain open, Hermite, Lindemann and Weierstrass proved in their seminal works the transcendence of $e$ and $\pi$ over a century ago [11, 12, 14, 22]. In fact, the celebrated Lindemann-Weierstrass theorem states that the exponential $e^{\alpha}$ is transcendental for every algebraic number $\alpha\neq 0$. Our main results imply that no similar statement holds in the context of (non-)normal numbers. Indeed, we construct a non-normal number $x$ such that all exponentials $e^{\alpha x}$ are not normal for every algebraic number $\alpha$.
Our main insight for the construction of non-normal numbers is to recall that despite being a set of Lebesgue measure zero, the non-normal form a co-meager set and are thus topologically big [2, 16, 17]. In this work, we introduce the closely related topological concept of super-density which is of independent interest and will allow us to prove non-normal numbers with desired properties. First, let us recall the definition of nowhere constant functions:
@@
We remark that for a function $\varphi$ to be nowhere constant on some interval $I$, it suffices that the set
@@
is dense in $I$. In particular, a Brownian motion is almost surely (with respect to the canonical Wiener measure) a nowhere constant continuous functions. Our main novel concept is content of the following
@@
The definition of super-density seems to be not been formulated in the mathematical literature before. However, certain constructions are based on related ideas. For instance, super-density is vaguely related to the definition of a universally meager set in [21]. Clearly, a super-dense set is dense in the usual sense with respect to the Euclidean metric; and it is also immediate that a super-dense set is necessarily uncountable. Hence, super-dense sets need to be topologically big and the natural question arises to understand the relation between super-density and the well-studied Baire categories. We briefly recall that a real subset $A\subset{\mathbb{R}}$ is called meager or of first category if it is given by an at most countable union of nowhere dense sets. Otherwise, $A$ is said to be of second category or a fat / non-meager set. If the complement $A^{c}:={\mathbb{R}}\setminus A$ is meager, we call $A$ co-meager. Finally, a set $A$ has the Baire property if and only if there is an open set $U$ and a meager set $N$ such that $A=U\Delta N$, where $\Delta$ denotes the symmetric difference. Our main result below gives a satisfactory relation between super-density and the Baire categories.
@@
The following hold.
  1. 1.
    Suppose $A\subset\mathbb{R}$ has the Baire property. Then, $A$ is super-dense if and only if $A$ is co-meager.
  2. 2.
  3. 3.
The proof of Theorem 1.3 is spelled out in Section 2.1. A few remarks are in order.
  1. 1.
    For general sets $A$ we have the chain of implications $A$ co-meager $\Rightarrow$ AA super-dense $\Rightarrow$ AA non-meager. This follows from the observation that co-meager and meager sets always have the Baire property.
  2. 2.
    The set $A$ from the third assertion is an example of a super-dense set not being co-meager. In view of the first assertion this set $A$ is also an example for a super-dense set without the Baire property.
  3. 3.
    By the Baire category theorem, at least one of the sets $A$ or $A^{c}$ from the second assertion is non-meager and, hence, an example for a set of second category which is not super-dense.
  4. 4.
    Super-density is in general not stable under intersection, that is, there exist super-dense sets $A,B$ such that $A\cap B$ is not super-dense itself. This is immediate from the third assertion.
To formulate the implications of Theorem 1.3 for normal numbers, we recall first some standard terminology. A real number $x\in{\mathbb{R}}$ has a unique expansion with respect to an integer base $b\geq 2$ in the form $x=X_{1}X_{2}\ldots X_{K}.x_{1}x_{2}\cdots$ with digits $X_{i},x_{i}\in\{0,1,\ldots,b-1\}$. This $b$-ary expansion of $x$ is unique if we agree to use a finite expansion, i.e., an expansion with only finitely many nonzero digits, whenever possible. We denote by $\Lambda_{b,d,M}(x)$ the relative fraction of the digit $d$ among the first $M$ fractional $x_{i}$ in base $b$ (see also (2.3)). A real number $x$ is called simply $\mathit{b}$-normal if $\lim_{M\to\infty}\Lambda_{b,d,M}(x)=\frac{1}{b}$, that is if all digits are asymptotically equidistributed in the $b$-ary expansion of $x$. If not just the digits but all finite words $w\in\bigcup_{k=1}^{\infty}\{0,1,\ldots,b-1\}^{k}$ appear with asymptotic frequency $b^{-|w|}$ in the $b$-ary expansion of $x$, we call $x$ a $\mathit{b}$-normal number. Equivalently, a real number $x$ is $b$-normal if and only if $x$ is simultaneously simply $b^{N}$-normal for all natural numbers $N\geq 1$. Finally, $x$ is an absolutely normal number or just a normal number if $x$ is $b$-normal with respect to all integer bases $b\geq 2$. Absolute normality is equivalent to the seemingly weaker condition that a number $x$ is simultaneously simply $b$-normal for all bases $b\geq 2$.
We denote by $\mathcal{N}_{b,s}=\{x\in{\mathbb{R}}\,|\,x\text{ is simply }b\text{-normal}\}$ the set of all $b$-simply normal number for some fixed integer base $b\geq 2$. Analogously, we introduce the set of all $b$-normal real numbers $\mathcal{N}_{b}$, and the set of all (absolutely) normal numbers $\mathcal{N}$. Finally, we introduce the set
$\mathcal{Z}_{b}:=\{x\in{\mathbb{R}}\,|\,\limsup_{M\to\infty}\Lambda_{b,0,M}(x)=1\},$ (1.3)
which contains all numbers which accumulate zeros in their $b$-adic expansion. Theorem 1.3 implies the following
Corollary 1.4.
The following hold.
  1. 1.
    The set of normal numbers $\mathcal{N}$ is not super-dense.
  2. 2.
    The sets of non-normal numbers $\mathcal{N}^{c}$ and $\mathcal{Z}_{b}$ are super-dense for each integer base $b\geq 2$.
  3. 3.
    $\mathcal{Z}_{b}$ has zero Hausdorff dimension for each integer base $b\geq 2$, whereas $\mathcal{N}^{c}$ has full Hausdorff dimension of $1$.
Proof.
That $\mathcal{N}^{c}$ has Hausdorff dimension $1$ is well known [10] and the proof that $\mathcal{Z}_{b}$ has zero Hausdorff dimension is presented in the appendix in form of Lemma A.1. In view of the first statement Theorem 1.3, it is enough to show that $\mathcal{Z}_{b}$ is co-meager since $\mathcal{N}^{c}$ is then co-meager and super-dense as larger set; and $\mathcal{N}$ is meager, hence not super-dense. The following argument is folklore and presented to be self-contained - refined and more subtle analysis can be found in [2, 16, 17].
We fix an integer base $b\geq 2$ and for $x\in{\mathbb{R}}$ we denote by $x_{i}\in\{0,\ldots,b-1\}$ the decimal places of $x$. We define
$G_{b}:=\bigcap_{n\geq 1}G_{n}$ (1.4)
We see that super-density and the measure-theoretical size of a set are independent of each other. Corollary 1.4 says: There is no nowhere constant continuous function that maps all non-normal numbers to normal numbers. However, since the proof of Theorem 1.3 is non-constructive, we do not obtain a procedure to compute a non-normal $x$ such that $\varphi(x)$ is also non-normal for a given nowhere constant continuous function $\varphi$. The next step - the construction of simultaneously non-normal numbers - is addressed in our next result.
@@
Given a non-empty open interval $I\subset{\mathbb{R}}$, a nowhere constant continuous function $\varphi\colon I\to{\mathbb{R}}$ and integer base $b\geq 2$, there exists an iterative procedure which computes a real number $x\in I$ such that
$x\in\mathcal{Z}_{b}\,\text{ and }\,\varphi(x)\in\mathcal{Z}_{b}$ (1.5)
In fact, we have more generally an iterative procedure, which, given a countable collection $(\varphi_{k})_{k\in{\mathbb{N}}}$ of nowhere continuous functions $\varphi_{k}\colon I\to{\mathbb{R}}$, determines a real number $x\in I$ such that
$\varphi_{k}(x)\in\mathcal{Z}_{b}\text{ for all }k\in\mathcal{Z}_{b}$ (1.6)
We give a few comments on Theorem 1.5.
@@
    The proof of Theorem 1.5 is spelled out in Section 2.2. Our method is constructive and the output is computable in the sense that under a few regularity assumptions on the functions $\varphi_{k}$ we can infer from the proof an algorithm computing inductively the digits of a real number $x$ such that all numbers $\varphi_{k}(x)$ are simultaneously non-normal. The details are presented in Corollary 2.9 below.
@@
    Theorem 1.5 gives a (nearly) optimal result on joint non-normality, which complements the recent work [15], where the construction of simultaneously normal numbers is addressed. In the future, we hope to address simultaneous normality and non-normality, that is, to construct for example a normal number $x$ such that $x^{2}$ is not normal - mimicking the conjectured property of $\sqrt{2}$.
@@
    Let us denote by $\bar{\mathbb{Q}}$ the set of real algebraic numbers. Applying Theorem 1.5 to the countable collection $(e^{\alpha x})_{\alpha\in\bar{\mathbb{Q}}\setminus\{0\}}$ and to all open intervals $I=(\alpha,\beta)\subset{\mathbb{R}}$, one concludes that the set $A:=\{x\,|\,e^{\alpha x}\text{ not normal for all }\alpha\in\bar{\mathbb{Q}}\setminus\{0\}\}$ forms a dense subset of $\mathbb{R}$}. This shows that as announced before that there is no analog of the Lindemann-Weierstrass theorem for normal numbers. As far as we know, this work is the first to give a constructive method to generate a non-normal number $x\neq 0$ for which $e^{x}$ is not normal.
@@
    Equation 1.6 implies
    $\bigcap_{k=1}^{\infty}\varphi_{k}(I\cap\mathcal{N}_{b,s}^{c})\neq\emptyset,$ (1.7)
    which sharpens the super-density of the set $\mathcal{N}_{b,s}^{c}$.
  5. 5.
    In the setting of a single nowhere constant function $\varphi$, one can also derive the super-density of the “absolutely non-normal” numbers $\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}$. The proof is presented in the appendix as Corollary B.1.
  6. 6.
    We conjecture that there is in fact no non-constant continuous function $\varphi$ such that $\varphi(\mathcal{N}^{c})\cap\mathcal{N}^{c}=\emptyset$. However, our constructive proof fails in this situation and it is unclear to us if the statement of Theorem 1.5 still holds for non-constant (but not necessarily nowhere constant) functions. We hope to address this question in future research.
The normal numbers $\mathcal{N}$ form the archetype of a set of full measure which is topologically small, i.e., meager and hence not super-dense. It is thus of interest to understand if in this case a map $\varphi$ contradicting super-density relies on nonconstructive tools. It turns out that we are able to explicitly construct a nowhere constant continuous function with the desired properties.
Theorem 1.6.
There is an explicit nowhere constant continuous function $\widehat{C}$$\mathbb{R}$$\to{\mathbb{R}}$ such that
$\widehat{C}(\mathcal{N})\cap\mathcal{N}=\emptyset$ (1.8)
Moreover, the function $\widehat{C}$$\mathbb{R}$$\to{\mathbb{R}}$ is computable.
The function $\widehat{C}$ consists of infinitely many self-similar copies of the well-known Cantor function. The explicit definition and a proof of the desired properties will be given in Section 2.3, where we also discuss the precise definition of computablility.
@@ 2 Proof of Main Results
In this section, we prove all statements of Theorem 1.3 one by one. The proofs are all independent of each other and are therefore presented separately. The proof of the first assertion requires several preparatory result. We start by showing that super dense sets cannot be meager.